ohyhyh:

The graph below shows the half life values of parent isotopes. The bar graph shows 'Parent Isotope' labeled on the x axis and 'Half-life (in billion years)' labeled on the y axis. The first bar is labeled Uranium-238 and its height is 4. The second bar is labeled 'Thorium-235' and its height is 14, the third bar is labeled 'Potassium-40' and its height is 1.5', and the fourth bar is labeled 'Rubidium-87' and its height is 49. Based on the graph, it can be concluded that the isotope which is most likely to be found with 6.25% of its original amount in 16 billion years? Rubidium-87 Uranium-238 Potassium-40 Thorium-235

ohyhyh:

1 attachment
KyanTheDoodle:

Ok, I'm not gonna try to plug a formula for this one. Putting this simply, every substance starts at 100% of it's value. After a half life, it will be 50%. Another half life, it will be 25%. How many half lifes will it take to get to 6.25%?

ohyhyh:

5

KyanTheDoodle:

100/2 = 50 50/2 = 25 25/2 = 12.5 12.5/2 = 6.25

ohyhyh:

so 4

KyanTheDoodle:

yes

ohyhyh:

so what next?

KyanTheDoodle:

so now that we know how many half lives it takes to get to 6.25%, let's divide the given time, 16 billion years, by 4

KyanTheDoodle:

and that will get you the half life of the right element

ohyhyh:

c?

KyanTheDoodle:

potassium has a half life of 1.5 billion, it would go through over 10 half lives by then

ohyhyh:

d?

KyanTheDoodle:

what is 16 divided by 4?

ohyhyh:

4

KyanTheDoodle:

which element has a half life of 4 billion years?

ohyhyh:

i mean itd have to be one of the two left but idk bc im confused

ohyhyh:

again

KyanTheDoodle:

let's start from the beginning

KyanTheDoodle:

how many half lives does it take to get to 6.25% of the original size?

ohyhyh:

4

KyanTheDoodle:

so then in 16 billion years, the element we're looking for would have to go through 4 half lives

KyanTheDoodle:

which means in one half life, it would be 16 billion years divided by 4

KyanTheDoodle:

you said 16 billion divided by 4 is 4 billion years. which bar is at the value of 4 billion years?

ohyhyh:

b

KyanTheDoodle:

that's correct

ohyhyh:

lol only took me 3 tries lol

ohyhyh:

All scientific theories on the solar system are based on biased scientific data non-testable facts non-observable information study of current evidences

ohyhyh:

i think this one is d?

KyanTheDoodle:

yes

ohyhyh:

ok cool

ohyhyh:

finally on q30

ohyhyh:

Scientists have discovered a solar system at a distance of 2000 AU from Earth. What is the distance, in kilometers, between the new solar system and Earth? 300 million kilometers 300 billion kilometers 20 million kilometers 20 billion kilometers

ohyhyh:

b?

KyanTheDoodle:

every AU is 149,597,870.7 km

KyanTheDoodle:

oh wait

KyanTheDoodle:

sorry I didn't see you already had an answer

KyanTheDoodle:

yes that's correct

ohyhyh:

which of these statements best compares P waves and S waves of an earthquake? Both S waves and P waves travel through oxygen gas. Both S waves and P waves travel through carbon dioxide gas. S waves do not tavel through molten rocks but P waves travel through molten rocks. P waves do not tavel through molten rocks but S waves travel through molten rocks.

ohyhyh:

i think c

KyanTheDoodle:

A and B are comparisons, C and D are contrasts

KyanTheDoodle:

but I'm starting to think whoever wrote the question didn't care about that

ohyhyh:

yeah

KyanTheDoodle:

yeah C is right

ohyhyh:

okay cool

ohyhyh:

im gonna start a new question